OpenStudy (anonymous):

what is x = to? 5x+2+6x-4=2x-25

OpenStudy (anonymous):

5x + 2 + 6x - 4 = 2x - 25 Well, combine like terms first so it would be: 11x -2 = 2x + 25 Then Add 2 to each side 11x = 2x + 27 Minus 2x from each side: 9x = 27 Divde each side by 9: 9/9 = 27/9 x=3

OpenStudy (anonymous):

Thank you (:

OpenStudy (anonymous):

No Problem! Hope that helped you understand it!

OpenStudy (anonymous):

yeah it did, for some reason i was getting 2.5